OpenStudy (lauluisa):

Write an equation of the line in standard form with the given slope through the given point. slope = -3, (-4, -7) A. y-3x = 12 B. y+3x = -12 C. y+3x = -19 D. y-3x= 19

OpenStudy (anonymous):

start with \[y+7=-3(x+4)\] in the "point slope" form and then change it to standard form

OpenStudy (lauluisa):

Why 7? I don't get it..

OpenStudy (mayankdevnani):

hope you understand ! @lauluisa

OpenStudy (lauluisa):

Thanks! I just read it... so it gives me y+7 = -3x + 4 y = -3x +11 @mayankdevnani

OpenStudy (mayankdevnani):

\[\large \bf y+7=-3(x+4)\]

OpenStudy (mayankdevnani):

not y+7=-3x+4

OpenStudy (lauluisa):

Yes I know that, but I have to solve... Right?

OpenStudy (mayankdevnani):

right

OpenStudy (lauluisa):

Ok, so that's what it gives me... y = -3x +11

OpenStudy (mayankdevnani):

\[\large \bf y+7=-3(x+4)\] \[\large \bf y+7=-3x-12\] \[\large \bf y+3x=-19=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

you made a mistake ! @lauluisa

OpenStudy (lauluisa):

Oh yes you right, I get it now lol Thanks!

OpenStudy (mayankdevnani):

don't apply this :- `-3x+4` its `-3(x+4)`

OpenStudy (lauluisa):

I forgot to multiply 3 times 4...

OpenStudy (lauluisa):

-3*

OpenStudy (mayankdevnani):

now understood ??

OpenStudy (anonymous):

that damned distributive property gets you every time steps are always the same first DISTRIBUTE

OpenStudy (lauluisa):

Yes thank you.. @mayankdevnani

OpenStudy (mayankdevnani):

good !!

OpenStudy (mayankdevnani):

and welcome :)